# QDepth-VLA: Quantized Depth Prediction as Auxiliary Supervision for Vision–Language–Action Models

### v1

@@ 4.1.1. Training Recipe
The QDepth-VLA based on open $\pi_{0}$ Ren (2024) is initially pre-trained for 9 epochs on the Fractal dataset Brohan et al. (2022), followed by 20 epochs of pre-training on the LIBERO-90 dataset Liu et al. (2023). After pre-training, the model is further fine-tuned on the four LIBERO subsets — Spatial, Object, Goal, and Long Liu et al. (2023) for around 50 epochs. For the Simpler benchmark Li et al. (2024a), the model is instead trained from scratch, first using the Bridge dataset Walke et al. (2023) for 13 epochs, and then the Fractal dataset Brohan et al. (2022) for an additional 9 epochs.
@@ 4.1.2. Evaluation Setup
For evaluation on LIBERO Liu et al. (2023), we adopt its four benchmark suites (Spatial, Object, Goal, and Long). Following the preprocessing method in Kim et al. (2024), image resolution is first normalized to $256\times 256$ and then resized to $224\times 224$ as model input. We also apply a 180-degree rotation to all images and use only the main-view RGB observations. Each task is evaluated over 50 rollouts, with the average success rate reported.
On the Simpler benchmark Li et al. (2024a), the evaluation covers two distinct settings: (1) models trained on the Bridge dataset Walke et al. (2023) are tested on tasks involving the WidowX250 robot, and (2) models trained on the Fractal dataset Brohan et al. (2022) are tested on tasks for the Google Robot. We adopt the visual matching configuration from Simpler Li et al. (2024a), evaluating each task across multiple initial positions with 10 rollouts per configuration. Consequently, the total number of evaluations per task ranges from 240 to 2400.
@@ 4.1.3. Main Results
As shown in Table 2, QDepth-VLA consistently outperforms single-view baselines across the LIBERO suites. It achieves stronger performance on both fine-grained and long-horizon tasks, reaching 94.0% on the Goal tasks and 72.6% on the Long tasks, surpassing the single-view baseline CoT-VLA Zhao et al. (2025) by 6.4% and 3.6%, respectively. Compared with open $\pi_{0}$ Ren (2024), QDepth-VLA shows consistent improvements across all four subsets (Spatial, Object, Goal, and Long), with the largest gain of 8.8% observed on the Spatial tasks.
While QDepth-VLA operates with only a single RGB observation, its average success rate remains competitive with multi-view VLAs. Specifically, QDepth-VLA achieves a mean success rate only 0.1% lower than $\pi_{0}$-FAST Ren (2024), while exceeding 4D-VLA Zhang et al. (2025a) by 3.1% and DreamVLA Zhang et al. (2025b) by 4.5% on the Goal tasks. Moreover, it surpasses $\pi_{0}$-FAST Black et al. (2024) by 12.4% on the more challenging Long tasks. Although leading multi-view models such as 3D-CAVLA Bhat et al. (2025), GeoVLA Sun et al. (2025) and UniVLA Bu et al. (2025) achieve higher overall results, our experimental results demonstrate that depth-augmented supervision effectively compensates for the lack of multi-view observations and brings single-view VLAs closer to multi-view performance levels.
By extension, we further implement a multi-view variant of QDepth-VLA while maintaining the same setting that predicts latent depth tokens corresponding only to the current main-view image. As shown in Table 2, the multi-view QDepth-VLA consistently outperforms single-view baselines. It achieves an average success rate of 94.9%, surpassing DreamVLA by 0.1% and $\pi_{0}$ by 0.8% on the Spatial tasks, and reaches 90.0% success on the Long tasks. While 3D-CAVLA Bhat et al. (2025) and GeoVLA Sun et al. (2025) achieve higher average success rates, they require explicit point cloud or depth map inputs during inference — modalities that QDepth-VLA does not rely on.
@@
RT-2-X Brohan et al. (2023) 78.7 77.9 25.0 - 60.7
Octo-Base Team et al. (2024) 17.0 4.2 22.7 - 16.8
OpenVLA Kim et al. (2024) 16.3 46.2 35.6 - 27.7
RoboVLM finetuned Li et al. (2024b) 77.3 61.7 43.5 - 63.4
SpatialVLA finetuned Qu et al. (2025) 86.0 77.9 57.4 - 75.1
Open $\pi_{0}$ Ren (2024) 97.5 87.1 68.0 32.9 71.4
@@
Octo-Base Team et al. (2024) 8.3 43.1 12.5 0.0 16.0
OpenVLA Kim et al. (2024) 0.0 4.1 0.0 0.0 1.0
RoboVLM finetuned Li et al. (2024b) 25.0 58.3 29.2 12.5 31.3
SpatialVLA finetuned Qu et al. (2025) 25.0 100.0 16.7 29.2 42.7
Open $\pi_{0}$ Ren (2024) 61.3 89.6 73.7 15.8 60.0
@@
Tables 3 and 4 present the experimental results of QDepth-VLA in the Simpler Li et al. (2024a) simulation environment.
As shown in Table 3, QDepth-VLA achieves a success rate of 98.3% on the pick coke can task, surpassing open $\pi_{0}$ Ren (2024) by 0.8% and SpatialVLA Qu et al. (2025) by 12.3%. On the more complex open top drawer and put apple in task, QDepth-VLA also attains 62.6%, outperforming open $\pi_{0}$ Ren (2024) by a large margin of 29.7%. This substantial improvement on long-horizon tasks can be attributed to enhanced spatial perception and object localization provided by depth-guided supervision, which improves the model’s ability to accurately identify and grasp target objects. As a result, the success probability of intermediate manipulation steps—such as grasping or placing—is increased, leading to higher overall task completion rates.
In Table 4, QDepth-VLA consistently achieves high success rates across various manipulation tasks. Notably, on the stack block task—which demands precise spatial reasoning and fine-grained control—QDepth-VLA reaches a success rate of 39.6%, surpassing SpatialVLA Qu et al. (2025) by 10.4%. Furthermore, on the Put Eggplant in Basket and Put Spoon on Towel tasks, QDepth-VLA achieves 95.0% and 82.0% success rates, respectively, outperforming open $\pi_{0}$ Ren (2024) by 5.4% and 8.3%. These improvements highlight the effectiveness of quantized depth supervision in enhancing spatial reasoning and manipulation precision, particularly for tasks involving object placement and coordination in cluttered 3D environments.
@@
To further validate the effectiveness of our depth supervision, we visualize depth reconstructions by passing the quantized predicted features through a trained VQ-VAE van den Oord et al. (2017) decoder. As shown in Fig. 3, the reconstructions preserve structural details and align well with object boundaries, demonstrating that the learned depth representations capture spatial geometry in a meaningful way.
@@ 4.2. Real-Robot Experiments
ACT Zhao et al. (2023) 20.0 0.0 0.0 0.0 5.0
Open $\pi_{0}$ Ren (2024) 50.0 40.0 40.0 0.0 32.5
@@ 4.2.2. Main Results
We evaluate QDepth-VLA on a series of pick-and-place tasks with varying difficulty to assess its spatial perception and localization ability. We also compare our method against representative baselines such as ACT Zhao et al. (2023). As shown in Table 5, QDepth-VLA consistently outperforms ACT Zhao et al. (2023) across all tasks, while ACT Zhao et al. (2023) fails to perform reliably in these complex real-world environments, QDepth-VLA achieves robust success rates. Compared to our baseline open $\pi_{0}$ Ren (2024), QDepth-VLA achieves a 20.0% improvement on the simple task of picking a banana, and further achieves gains of 10.0% on both Task 3 and Task 4, demonstrating stronger generalization to challenging scenarios and tasks.
@@ 4.3. Ablation Study
To evaluate the contribution of each proposed component, we perform a series of controlled ablation experiments in the Simpler Li et al. (2024a) simulation environment. Four ablated variants are considered: (1) removing the depth supervision signal by setting its loss weight to zero (w/o Depth Loss); (2) removing the dedicated depth prediction branch (w/o Depth Expert); (3) replacing latent depth token prediction with pixel-wise regression (w/o Pixel Prediction); and (4) substituting the proposed hybrid attention mask with a standard version that enforces proprioception-to-depth attention (w/o Hybrid Attn). Quantitative results are reported in Table 6.
@@
In this variant, the proposed hybrid attention mask is replaced with a DreamVLA-style Zhang et al. (2025b) configuration, removing dynamic and semantic modalities. This setting tests whether relative depth maps can enhance proprioceptive state perception and thereby improve action generation quality. As expected, the performance declines by 5.5% on average, with the most substantial drop on the Carrot Task (-15.8%). This result indicates that enforcing proprioception-to-depth attention introduces noise rather than useful guidance, as relative depth lacks absolute positional encoding necessary for stable control.

### v2

@@ 4.1.1. Training Recipe
The QDepth-VLA based on open $\pi_{0}$ Ren2024OpenPiZero is initially pre-trained for 9 epochs on the Fractal dataset Brohan2022RT1RT, followed by 20 epochs of pre-training on the LIBERO-90 dataset Liu2023LIBEROBK. After pre-training, the model is further fine-tuned on the four LIBERO subsets — Spatial, Object, Goal, and Long Liu2023LIBEROBK for around 50 epochs. For the Simpler benchmark Li2024EvaluatingRR, the model is instead trained from scratch, first using the Bridge dataset Walke2023BridgeDataVA for 13 epochs, and then the Fractal dataset Brohan2022RT1RT for an additional 9 epochs.
@@ 4.1.2. Evaluation Setup
For evaluation on LIBERO Liu2023LIBEROBK, we adopt its four benchmark suites (Spatial, Object, Goal, and Long). Following the preprocessing method in Kim2024OpenVLAAO, image resolution is first normalized to $256\times 256$ and then resized to $224\times 224$ as model input. We also apply a 180-degree rotation to all images and use only the main-view RGB observations. Each task is evaluated over 50 rollouts, with the average success rate reported.
On the Simpler benchmark Li2024EvaluatingRR, the evaluation covers two distinct settings: (1) models trained on the Bridge dataset Walke2023BridgeDataVA are tested on tasks involving the WidowX250 robot, and (2) models trained on the Fractal dataset Brohan2022RT1RT are tested on tasks for the Google Robot. We adopt the visual matching configuration from Simpler Li2024EvaluatingRR, evaluating each task across multiple initial positions with 10 rollouts per configuration. Consequently, the total number of evaluations per task ranges from 240 to 2400.
@@ 4.1.3. Main Results
As shown in Table 2, QDepth-VLA consistently outperforms single-view baselines across the LIBERO suites. It achieves stronger performance on both fine-grained and long-horizon tasks, reaching 94.0% on the Goal tasks and 72.6% on the Long tasks, surpassing the single-view baseline CoT-VLA Zhao2025CoTVLAVC by 6.4% and 3.6%, respectively. Compared with open $\pi_{0}$ Ren2024OpenPiZero, QDepth-VLA shows consistent improvements across all four subsets (Spatial, Object, Goal, and Long), with the largest gain of 8.8% observed on the Spatial tasks.
While QDepth-VLA operates with only a single RGB observation, its average success rate remains competitive with multi-view VLAs. Specifically, QDepth-VLA achieves a mean success rate only 0.1% lower than $\pi_{0}$-FAST Ren2024OpenPiZero, while exceeding 4D-VLA Zhang20254DVLASV by 3.1% and DreamVLA Zhang2025DreamVLAAV by 4.5% on the Goal tasks. Moreover, it surpasses $\pi_{0}$-FAST Black20240AV by 12.4% on the more challenging Long tasks. Although leading multi-view models such as 3D-CAVLA Bhat20253DCL, GeoVLA Sun2025GeoVLAE3 and UniVLA Bu2025UniVLALT achieve higher overall results, our experimental results demonstrate that depth-augmented supervision effectively compensates for the lack of multi-view observations and brings single-view VLAs closer to multi-view performance levels.
By extension, we further implement a multi-view variant of QDepth-VLA while maintaining the same setting that predicts latent depth tokens corresponding only to the current main-view image. As shown in Table 2, the multi-view QDepth-VLA consistently outperforms single-view baselines. It achieves an average success rate of 94.9%, surpassing DreamVLA by 0.1% and $\pi_{0}$ by 0.8% on the Spatial tasks, and reaches 90.0% success on the Long tasks. While 3D-CAVLA Bhat20253DCL and GeoVLA Sun2025GeoVLAE3 achieve higher average success rates, they require explicit point cloud or depth map inputs during inference — modalities that QDepth-VLA does not rely on.
@@
RT-2-X Brohan2023RT2VM 78.7 77.9 25.0 - 60.7
Octo-Base Team2024OctoAO 17.0 4.2 22.7 - 16.8
OpenVLA Kim2024OpenVLAAO 16.3 46.2 35.6 - 27.7
RoboVLM finetuned Li2024TowardsGR 77.3 61.7 43.5 - 63.4
SpatialVLA finetuned Qu2025SpatialVLAES 86.0 77.9 57.4 - 75.1
Open $\pi_{0}$ Ren2024OpenPiZero 97.5 87.1 68.0 32.9 71.4
@@
Octo-Base Team2024OctoAO 8.3 43.1 12.5 0.0 16.0
OpenVLA Kim2024OpenVLAAO 0.0 4.1 0.0 0.0 1.0
RoboVLM finetuned Li2024TowardsGR 25.0 58.3 29.2 12.5 31.3
SpatialVLA finetuned Qu2025SpatialVLAES 25.0 100.0 16.7 29.2 42.7
Open $\pi_{0}$ Ren2024OpenPiZero 61.3 89.6 73.7 15.8 60.0
@@
Tables 3 and 4 present the experimental results of QDepth-VLA in the Simpler Li2024EvaluatingRR simulation environment.
As shown in Table 3, QDepth-VLA achieves a success rate of 98.3% on the pick coke can task, surpassing open $\pi_{0}$ Ren2024OpenPiZero by 0.8% and SpatialVLA Qu2025SpatialVLAES by 12.3%. On the more complex open top drawer and put apple in task, QDepth-VLA also attains 62.6%, outperforming open $\pi_{0}$ Ren2024OpenPiZero by a large margin of 29.7%. This substantial improvement on long-horizon tasks can be attributed to enhanced spatial perception and object localization provided by depth-guided supervision, which improves the model’s ability to accurately identify and grasp target objects. As a result, the success probability of intermediate manipulation steps—such as grasping or placing—is increased, leading to higher overall task completion rates.
In Table 4, QDepth-VLA consistently achieves high success rates across various manipulation tasks. Notably, on the stack block task—which demands precise spatial reasoning and fine-grained control—QDepth-VLA reaches a success rate of 39.6%, surpassing SpatialVLA Qu2025SpatialVLAES by 10.4%. Furthermore, on the Put Eggplant in Basket and Put Spoon on Towel tasks, QDepth-VLA achieves 95.0% and 82.0% success rates, respectively, outperforming open $\pi_{0}$ Ren2024OpenPiZero by 5.4% and 8.3%. These improvements highlight the effectiveness of quantized depth supervision in enhancing spatial reasoning and manipulation precision, particularly for tasks involving object placement and coordination in cluttered 3D environments.
@@
To further validate the effectiveness of our depth supervision, we visualize depth reconstructions by passing the quantized predicted features through a trained VQ-VAE Oord2017NeuralDR decoder. As shown in Fig. 3, the reconstructions preserve structural details and align well with object boundaries, demonstrating that the learned depth representations capture spatial geometry in a meaningful way.
@@ 4.2. Real-Robot Experiments
ACT Zhao2023LearningFB 20.0 0.0 0.0 0.0 5.0
Open $\pi_{0}$ Ren2024OpenPiZero 50.0 40.0 40.0 0.0 32.5
@@ 4.2.2. Main Results
We evaluate QDepth-VLA on a series of pick-and-place tasks with varying difficulty to assess its spatial perception and localization ability. We also compare our method against representative baselines such as ACT Zhao2023LearningFB. As shown in Table 5, QDepth-VLA consistently outperforms ACT Zhao2023LearningFB across all tasks, while ACT Zhao2023LearningFB fails to perform reliably in these complex real-world environments, QDepth-VLA achieves robust success rates. Compared to our baseline open $\pi_{0}$ Ren2024OpenPiZero, QDepth-VLA achieves a 20.0% improvement on the simple task of picking a banana, and further achieves gains of 10.0% on both Task 3 and Task 4, demonstrating stronger generalization to challenging scenarios and tasks.
@@ 4.3. Ablation Study
To evaluate the contribution of each proposed component, we perform a series of controlled ablation experiments in the Simpler Li2024EvaluatingRR simulation environment. Four ablated variants are considered: (1) removing the depth supervision signal by setting its loss weight to zero (w/o Depth Loss); (2) removing the dedicated depth prediction branch (w/o Depth Expert); (3) replacing latent depth token prediction with pixel-wise regression (w/o Pixel Prediction); and (4) substituting the proposed hybrid attention mask with a standard version that enforces proprioception-to-depth attention (w/o Hybrid Attn). Quantitative results are reported in Table 6.
@@
In this variant, the proposed hybrid attention mask is replaced with a DreamVLA-style Zhang2025DreamVLAAV configuration, removing dynamic and semantic modalities. This setting tests whether relative depth maps can enhance proprioceptive state perception and thereby improve action generation quality. As expected, the performance declines by 5.5% on average, with the most substantial drop on the Carrot Task (-15.8%). This result indicates that enforcing proprioception-to-depth attention introduces noise rather than useful guidance, as relative depth lacks absolute positional encoding necessary for stable control.